A large container of water is placed on a set of scales which initially read 0.500 kg. The container has a overflow spout that can be used to collect any overflow into a second container.

The second container is also on a set of scales and is initially empty. The second set of scales initially read 0.065 kg.

A small ball is lowered into the large container, displacing water down the overflow spout and into the second container.

The readings on both scales are now 0.585 kg and 0.100 kg respectively.

What is the density of the ball? (to 2 s.f and in kg m−3)

(ρwater = 1000 kg m−3)


1
Expert's answer
2021-01-23T14:07:05-0500
"V_b=\frac{0.1-0.065}{1000}=0.000035\ m^3"

"m_b=0.585-0.5+(0.1-0.065)=0.12\ kg\\
\rho_b=\frac{0.12}{0.000035}=3400\frac{kg}{m^3}"